OpenStudy (anonymous):

please divide x^2-x-14 by x-4

OpenStudy (lgbasallote):

first rewrite it as ______________ x - 4 | x^2 - x - 14 now...what is x^2 divided by x?

OpenStudy (anonymous):

x

OpenStudy (lgbasallote):

good x ______________ x - 4 | x^2 - x - 14 now we multiply x to each term of x - 4 x ______________ x - 4 | x^2 - x - 14 x^2 - 4x --------------- we subtract can you subtract x^2 - x - (x^2 - 4x)?

OpenStudy (anonymous):

yes would it simplify to -5x?

OpenStudy (lgbasallote):

-5x? hmm remember you distribute the minus sign... x^2 - x - x^2 - (-4x) it becomes x^2 - x - x^2 + 4x

OpenStudy (anonymous):

ok got it so then you would subtract x from 4x?

OpenStudy (lgbasallote):

yup!!

OpenStudy (lgbasallote):

so what do you have?

OpenStudy (anonymous):

I believe 3x because the x^2s cancel out ?

OpenStudy (lgbasallote):

you got it! x ______________ x - 4 | x^2 - x - 14 x^2 - 4x --------------- 3x now we bring down 14 x ______________ x - 4 | x^2 - x - 14 x^2 - 4x --------------- 3x - 14 then we repeat the process...divide 3x by x

OpenStudy (anonymous):

3?

OpenStudy (lgbasallote):

good.. x + 3 ______________ x - 4 | x^2 - x - 14 x^2 - 4x --------------- 3x - 14 now multiply 3 to each term of x - 4

OpenStudy (anonymous):

would it be 3x-12?

OpenStudy (lgbasallote):

yup!

OpenStudy (lgbasallote):

x + 3 ______________ x - 4 | x^2 - x - 14 x^2 - 4x --------------- 3x - 14 3x - 12 -------- subtract again

OpenStudy (anonymous):

thank you so much, could you help me with solving rad72-rad50?

OpenStudy (lgbasallote):

do you know the rest for that one?

OpenStudy (lgbasallote):

remember to put the remainder :D

OpenStudy (anonymous):

do you mean the division problem is not finished

OpenStudy (lgbasallote):

haha yeah :))

OpenStudy (lgbasallote):

you still have to subtract 3x - 12 from 3x - 14

OpenStudy (anonymous):

so the final answer is -2?

OpenStudy (lgbasallote):

we still have to write -2 as a remainder to do that we write \[\frac{remainder}{divisor}\] since our remainder is -2 and our divisor is 2x + 1 it becomse \[\frac{-2}{2x+1}\] got it? we add that to the quotient we have so it's \[\large x + 3 - \frac{2}{2x+1}\]

OpenStudy (lgbasallote):

for more info on dividing polynomials by binomials check this out (a tutorial i made myself :D) http://openstudy.com/updates/4f9c9dace4b000ae9ed1966f now...as for your second question...please post in another question because long threads tend to be laggy

OpenStudy (anonymous):

ok thanks